How do you seperate 116 ito two parts,where one of which shall be 3 times the other?
Answer by Fermat(136)

let one part be x
let the other part be y
then x = 3y, say
Also,
116 = x + y
substitute for x = 3y into the above eqn
116 = 3y + y
116 = 4y
divide both sides by 4,
29 = y
y = 29
======
x = 87 (87 = 3*29)
======